The equilibrium constant, K, for the following reaction is 1.80×10-2 at 698 K.

2HI(g) H2(g) + I2(g)

An equilibrium mixture of the three gases in a 1.00 L flask at 698 K contains 0.319 M HI, 4.27×10-2 M H2 & 4.27×10-2 M I2.

What will be the concentrations of the three gases once equilibrium has been reestablished, if 0.224 mol of HI(g) is added to the flask?

[HI] = M

[H2] = M

[I2] = M

please help me!
Answer #1

From Le-Chatelier's Principle, If reaction at equilibrium is subjected to a change in any of the reaction condition like pressure, volume, concentration etc then reaction will shift in a direction to undo the effect of change.

Hence, if we add more HI to the beaker then reaction will shift in forward direction to decrease the added amount

Since volume = 1L

So, moles = molarity

So,

2HI (g) H2(g) I2(g)
Initial (M) 0.319 4.27 *10-2 = 0.0427 0.0427
added (M) 0.224 0 0
Change (M) -2x x +x
Equilibrium (M) 0.543 -2x 0.0427+x 0.0427+x

Now, Kc = [H2][I2]/[HI]2 = 1.80*10-2

1.80*10-2 = (0.0427+x)2 / (0.543-2x)2

0.134 = (0.0427+x) / (0.543-2x)

0.0729 -0.268x = 0.0427 +x

0.0302 = 1.268 x

x = 0.0238 M

Hence, [HI] = 0.543 -2(0.0238) = 0.495 M

[I2] = [H2] = 0.0427 + 0.0238 = 0.0665 M